Grade 11Physical Chemistry

the density of the substance at 1atm and 750k 0.3g/lit if the molecular weight of the substance is 27 then what are the forces existing among the gas molecules

Ashok
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
From,
zd = PM/RT
Putting al,, the values we get,
z = PM/dRT
z => 1*27/(0.3)*0.0821*750 = 1.46
Here compressibility factor that is z is greater than 1 hence the nature of forces would be repulsive.